OpenStudy (anonymous):

How many solutions are there to the quadratic equation 4x2 + 8x + 3 = 0? 0 1 2 none of the above

OpenStudy (anonymous):

stop spamming (again)

OpenStudy (anonymous):

And what did I tell you about looking for answers vs asking for help?

OpenStudy (anonymous):

well give me dat website again

OpenStudy (anonymous):

lol

OpenStudy (anonymous):

wolframalpha.com

OpenStudy (anonymous):

and its not spam its math questions lol

OpenStudy (anonymous):

It's spam when you post 5 in less than a minute.

OpenStudy (anonymous):

nope its called copy and paste lol

OpenStudy (anonymous):

i tld u my math class is an online class lol

OpenStudy (anonymous):

And I told you, that we're not the ones taking the class, you are.

OpenStudy (anonymous):

either learn the material, or stop wasting your time.

OpenStudy (anonymous):

well instead of writing stuff like dat teach the lesson to me instead of writng stuff that i already kno thank u very much

OpenStudy (anonymous):

Do you actually want to know how to do this?

OpenStudy (anonymous):

yes because i have to tell my teacher how i got the ansers and we do that by phone not online so i have to learn it

OpenStudy (anonymous):

Well you won't learn it by pasting the same question 5 times and having other people do it for you. You will learn it by posting one question and spending some time with someone who is willing to work through it with you.

OpenStudy (anonymous):

If you know how to do it you can solve them very quickly.

OpenStudy (anonymous):

its not the same question acytually its a series of questions lol

OpenStudy (anonymous):

It is the same question. Just because it has different numbers doesn't make it different.

OpenStudy (anonymous):

true dat ture dat lol well can u show me how then step by step??

OpenStudy (anonymous):

Do you know the quadratic formula?

OpenStudy (anonymous):

polpak are you like a lot older than davette

OpenStudy (anonymous):

I have no idea.

OpenStudy (anonymous):

lol

OpenStudy (anonymous):

okay then are you older then 18 is the q

OpenStudy (anonymous):

Do you know the quadratic formula Dave?

OpenStudy (anonymous):

nope can u show me ??

OpenStudy (anonymous):

and its davette im a girl just saying lol

OpenStudy (anonymous):

\[ax^2 + bx + c = 0 \implies x = {-b \pm \sqrt{b^2-4ac} \over 2a}\]

OpenStudy (anonymous):

So figure out what your a, b, and c is then plug them in to find x

OpenStudy (anonymous):

o ok

OpenStudy (anonymous):

do i do that for each problem??

OpenStudy (anonymous):

Any time you have a problem that looks like \[ax^2 + bx + c = 0\] (or you can rewrite in that form) You can solve it by pluggin in your a, b, and c to the quadratic equation.

OpenStudy (anonymous):

Davette and PolPak sitting in a tree .......